OpenStudy (anonymous):

Symmetry of the graph? F(x)= x^6-x^4+x^2-8

jimthompson5910 (jim_thompson5910):

what is f(-x) equal to?

OpenStudy (anonymous):

-f(x)

jimthompson5910 (jim_thompson5910):

f(x)= x^6-x^4+x^2-8 f(-x)= (-x)^6-(-x)^4+(-x)^2-8 ... replace every x with -x f(-x) = x^6-x^4+x^2-8

jimthompson5910 (jim_thompson5910):

So f(-x) = f(x) that means there is symmetry with respect to the y axis